In a survey, 30 people were asked how much they spent on their child's last birthday gift. The results were roughly bell-shaped with a mean of $33 and standard deviation of $10. Find the margin of error at a 99% confidence level.

Give your answer to two decimal places.
Answer #1

stution 7130 MAX - 33 SD = 10 4:0.01 z distribution (n>30) son : 30 50, Me we z distribution E = 2412 * 200 = 7.58 (99% confi